# Copula Conformal Prediction for Multi-step Time Series Forecasting

### 23 Jan 2024 v3

@@ Abstract
Accurate uncertainty measurement is a key step to building robust and reliable machine learning systems. Conformal prediction is a distribution-free uncertainty quantification algorithm popular for its ease of implementation, statistical coverage guarantees, and versatility for underlying forecasters. However, existing conformal prediction algorithms for time series are limited to single-step prediction without considering the temporal dependency. In this paper we propose the Copula Conformal Prediction algorithm for multivariate, multi-step Time Series forecasting, CopulaCPTS. We prove that CopulaCPTS has finite sample validity guarantee. On four synthetic and real-world multivariate time series datasets, we show that CopulaCPTS produces more calibrated and efficient confidence intervals for multi-step prediction tasks than existing techniques.
@@ 1 Introduction
Deep learning models are becoming widely used in high-risk settings such as healthcare and transportation. In these settings, it is important that a model produces calibrated uncertainty to reflect its own confidence. Confidence regions are a common approach to quantify prediction uncertainty (Khosravi et al., 2011). A $(1-\alpha)$-confidence region $\Gamma^{1-\alpha}$ for a random variable $y$ is valid if it contains $y$’s true value with high probability: ${\mathbb{P}}[y\in\Gamma^{1-\alpha}]\geq 1-\alpha.$ Note that one can make the confidence region infinitely large to satisfy validity. But for the confidence region to be useful, we also want to minimize its area while remaining valid; this is known as efficiency of the region.
Conformal prediction (CP) is a powerful method to produces confidence regions with finite-sample guarantees of validity (Vovk et al., 2005; Lei et al., 2018). Furthermore, it makes no assumptions about the prediction model or the underlying data distribution. CP’s generality, simplicity, and statistical guarantees have made it popular for many real-world applications including time series prediction (Xu & Xie, 2021), drug discovery (Eklund et al., 2015) and safe robotics (Luo et al., 2021).
This paper considers the setting of multi-step time series prediction in the presence of a training dataset of similar time series. We will motivate with an example application of car trajectory prediction, illustrated in figure 1. The task is to predicting future trajectory ${\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{k}$ given past trajectory ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{t}$. These time steps are temporally dependent. On the other hand, we also have a dataset of past trajectories ${\mathcal{D}}$, where the trajectories are independent of each other.
@@
There are many real-world task that presents the same challenges as our motivating example, such as EEG forecasting (each patient is independent), short-term weather forecasting (local meteorology history is independent), etc. They require predicting multiple time steps into the future, so it is desired to have a “cone of uncertainty” that covers the entire course of the forecasts. Existing CP methods for time series data either only provide coverage guarantee for individual time steps (Gibbs & Candes, 2021; Xu & Xie, 2021) or produce confidence regions often too inefficient to be useful, especially in long horizons or multivariate settings (Stankevičiūtė et al., 2021).
In this paper, we present a practical and effective conformal prediction algorithm for multi-step time series forecasting. We introduce CopulaCPTS, a Copula-based Conformal Prediction algorithm for multi-step Time Series forecasting. We improve efficiency by utilizing copulas to model the dependency between forecasted time steps. A copula is a multivariate cumulative distribution function that models the dependence between multiple random variables. By using copulas to model the uncertainty jointly over time, we can shrink the confidence regions significantly while maintaining validity. Copulas have been used for conformal prediction (Messoudi et al., 2021), but they focus on multiple target prediction in non-temporal settings and did not provide a validity proof.
@@
  • We introduce CopulaCPTS, a general uncertainty quantification algorithm that can be applied to any multivariate multi-step forecaster, with statistical guarantees of validity.
@@
  • CopulaCPTS can be extended to obtain valid confidence intervals for time series forecasts of varying lengths.
@@ 2 Related Work
Uncertainty Quantification for Deep Time-Series Forecasting.
The two major paradigms of Uncertainty Quantification (UQ) methods for deep neural networks are Bayesian and Frequentist. Bayesian approaches estimate a distribution over the model parameters given data, and then marginalize these parameters to form output distributions via Markov Chain Monte Carlo (MCMC) sampling (Welling & Teh, 2011; Neal, 2012; Chen et al., 2014) or variational inference (VI) (Graves, 2011; Kingma et al., 2015; Blundell et al., 2015; Louizos & Welling, 2017). Wang et al. (2019); Wu et al. (2021) propose Bayesian Neural Networks (BNN) for UQ of spatiotemporal forecasts. In practice, Bayesian UQ has two major drawbacks: (1) it requires significant modification to the training process of the models, and (2) BNNs can be computationally expensive and difficult to optimize, especially for larger networks (Lakshminarayanan et al., 2017; Zadrozny & Elkan, 2001). Therefore, UQ for deep neural network time series forecasts often adopts approximate Bayesian inference such as MC-dropout (Gal & Ghahramani, 2016b; Gal et al., 2017).
Frequentist UQ methods emphasize robustness against variations in the data. These approaches either rely on resampling the data or learning an interval bound to encompass the dataset. For time series forecasting UQ, frequentist approaches include ensemble methods such as bootstrap (Efron & Hastie, 2016; Alaa & Van Der Schaar, 2020) and jackknife methods (Kim et al., 2020; Alaa & Van Der Schaar, 2020); interval prediction methods include interval regression through proper scoring rules (Kivaranovic et al., 2020; Wu et al., 2021), and quantile regression (Takeuchi et al., 2006), with many recent advances for time series UQ (Tagasovska & Lopez-Paz, 2019; Gasthaus et al., 2019; Park et al., 2022; Kan et al., 2022). Many of the frequentist methods produce asymptotically valid confidence regions and can be categorized as distribution-free UQ techniques as they are (1) agnostic to the underlying model and (2) agnostic to data distribution.
@@
Conformal prediction (CP) is an important member of distribution-free UQ methods; we refer readers to Angelopoulos & Bates (2021) for a comprehensive introduction and survey. CP has become popular because of its simplicity, theoretical soundness, and low computational cost. A key difference between CP and other UQ methods is that under the exchangeability assumption, conformal methods guarantee validity in finite samples (Vovk et al., 2005).
Most relevant to our work is the recent endeavor in generalizing CP to time-series forecasting. According to Stankevičiūtė et al. (2021) there are two settings: data generated from (1) one single time series or (2) multiple independent time series. For the first setting, ACI (Gibbs & Candes, 2021) and EnbPI (Xu & Xie, 2021) developed CP algorithms that relax the exchangeability assumption while maintaining asymptotic validity via online learning (former) and ensembling (later); Zaffran et al. (2022) further improves online adaptation. Sousa et al. (2022) combines EnbPI with conformal quantile regression (Romano et al., 2019) to model heteroscedastic time series. However, because these algorithms operate on one single time series, the validity guarantees are on average over the predicted time steps and are asymptotic, rather than covering the full horizon as in our setting.
CopulaCPTS is developed for the setting where data is generated from many independent time series. Stankevičiūtė et al. (2021) shares the same setting but provides only a univariate solution. We show that their method of applying Bonferroni correction produces inefficient confidence regions, especially for multidimensional data or long prediction horizons. Messoudi et al. (2021) uses a copula function for multi-target CP for non-temporal data, creating box-like regions to account for the correlations between the labels. However, they do not provide a theoretical proof and empirical results have shown that are often invalid. This paper builds upon these works and presents a novel two-step algorithm with guaranteed multivariate multi-step coverage and efficient confidence regions.
@@ 3.1 Inductive Conformal Prediction (ICP)
${\mathbb{P}}(y\in\Gamma^{1-\alpha}(x))\geq 1-\alpha$ (1)
@@ 4 Copula Conformal Prediction for Time Series (CopulaCPTS)
Denote the time series dataset of size $n$ as $\mathcal{D}=\{z^{i}=(x_{1:t}^{i},y_{1:k}^{i})\}_{i=1}^{n}$, where $x_{1:t}\in\mathbb{R}^{t\times d}$ is $t$ input steps, and $y_{1:k}\in\mathbb{R}^{k\times d}$ is $k$ prediction steps, both with dimension $d$ at each step. Each data point $z^{i}$ is sampled i.i.d. from an unknown distribution $\mathcal{Z}$. In the multi-step forecasting setting, given a confidence level $1-\alpha$, the algorithm produces $k$ confidence regions for a test input $x_{1:t}^{n+1}$, denoted as $[\Gamma^{1-\alpha}_{1},\ldots,\Gamma^{1-\alpha}_{k}]$. We say the confidence regions are valid if all time steps in the forecast is covered:
@@ 4.1 Algorithm Details
The two calibration sets allows us to prove validity for both components of our algorithm. At the cost of using a subset of the data to calibrate a copula, our approach produces provably more efficient confidence regions compared to worst-case corrections such as union bounding in Stankevičiūtė et al. (2021) which is a lower bound for copulas (Appendix B.1), and more valid regions than Messoudi et al. (2021) (table 1).
@@
For brevity, we will use ${\mathcal{S}}1=\{s^{i}\}_{z^{i}\in\mathcal{D}_{cal-1}}$ to denote the set of nonconformity scores of data in $\mathcal{D}_{cal-1}$ and ${\mathcal{S}}2=\{s^{i}\}_{z^{i}\in\mathcal{D}_{cal-2}}$ the set of nonconformity scores of data in $\mathcal{D}_{cal-2}$. Subscript $j$ will be used to to index the set of specific time steps of the scores: ${\mathcal{S}}1_{j}=\{s^{i}_{j}\}_{z^{i}\in\mathcal{D}_{cal-1}}$, ${\mathcal{S}}2_{j}=\{s^{i}_{j}\}_{z^{i}\in\mathcal{D}_{cal-2}}$.
@@
We adopt the empirical copula (Ruschendorf, 1976) for modeling and proof in this work. The empirical copula is a non-parametric method of estimating marginals directly from observation, and hence does not introduce any bias. For the joint distribution of $k$ time steps, we construct $C_{\text{empirical}}:[0,1]^{k}\rightarrow[0,1]$ as eqn 7. Here boldface $\bm{\infty}$ is a k-dimensional vector with each $\bm{\infty}_{j}=\infty$ for $j=1,\ldots,k$.
@@
Note that the ${\mathbf{u}}^{*}$ is not and does not have to be unique; any solution that satisfy the constraint in Eq. 8 will guarantee multi-step validity (Appendix A). The minimization helps with efficiency, i.e. the sharpness of the confidence regions. We use a gradient descent algorithm for the optimization in implementation (see Appendix B.2 for details, and Appendix C.6 for a study of its effectiveness). Lastly, We obtain $(s_{1}^{*},\ldots,s_{k}^{*})$ by $\hat{F}_{j}^{-1}({\mathbf{u}}^{*}_{j})$ and construct the confidence region for each time step $j\in\{1,\ldots,k\}$ as the set of all $y_{j}\in\mathbb{R}^{d}$ such that the nonconformity score is less than $s_{j}^{*}$. Algorithm 1 summarizes the CoupulaCPTS procedure.
@@ 16       $\hat{y}^{i}_{1:k}\leftarrow\hat{f}(x_{1:t}^{i})$
Algorithm 1 Copula Conformal Time Series Prediction
@@ 5 Experiments
We compare our model with three representative works in different paradigms of uncertainty quantification for neural network time series prediction: the Bayesian-motivated Monte Carlo dropout RNN (MC-dropout) by Gal & Ghahramani (2016a), the frequentist blockwise jackknife RNN (BJRNN) by Alaa & Van Der Schaar (2020), a conformal forecasting RNN (CF-RNN) by Stankevičiūtė et al. (2021), and the multi-target copula algorithm that does not have the two step calibration (Copula) by Messoudi et al. (2021). We use the same underlying prediction model for post-hoc uncertainty quantification methods BJRNN, CF-RNN, and CopulaCPTS. The MC-dropout RNN is of the same architecture but is trained separately, as it requires an extra dropout step during training and inference.
@@ 5.2 Real world datasets
Comparison of models at different horizon lengths
@@
CopulaCPTS for Autoregressive prediction
@@ 6 Conclusion and Discussion
We also discuss the limitation of our algorithm. As CopulaCPTS require two calibration steps, it is suitable only when there are abundant data for calibration. The validity proof relies on using the empirical copula, so it does not apply to other learnable copula classes. Future work includes (1) improving the autoregressive extension of CopulaCPTS, to achieve coverage over the whole horizon, and (2) developing online settings onf CopulaCPTS for decision making.
@@
Let ${\mathcal{S}}1=\{s^{i}\}_{z^{i}\in\mathcal{D}_{cal-1}}$ be the set of nonconformity scores of data in $\mathcal{D}_{cal-1}$ and ${\mathcal{S}}2=\{s^{i}\}_{z^{i}\in\mathcal{D}_{cal-2}}$ the set of nonconformity scores of data in $\mathcal{D}_{cal-2}$. Subscript $j$ will be used to to index the set of specific time steps of the scores: ${\mathcal{S}}1_{j}=\{s^{i}_{j}\}_{z^{i}\in\mathcal{D}_{cal-1}}$, ${\mathcal{S}}2_{j}=\{s^{i}_{j}\}_{z^{i}\in\mathcal{D}_{cal-2}}$.
@@
Each ${\mathbf{u}}^{i}$ can be seen as a multivariate nonconformity score for data sample $z^{i}$. We will now illustrate that an empirical copula on $\mathcal{U}$ is a rank statistic, and hence we can apply the proof of conformal prediction to prove finite sample validity guarantee.
@@
Next, we define the empirical multivariate quantile function for $\mathcal{U}$, a set of $k$-dimensional vectors.
@@
  • The uncertain regions are constructed as as (Algorithm 1, line 17):

### 18 Mar 2024 v4

@@ Abstract
Accurate uncertainty measurement is a key step in building robust and reliable machine learning systems. Conformal prediction is a distribution-free uncertainty quantification framework popular for its ease of implementation, finite-sample coverage guarantees, and generality for underlying prediction algorithms. However, existing conformal prediction approaches for time series are limited to single-step prediction without considering the temporal dependency. In this paper, we propose the Copula Conformal Prediction algorithm for multivariate, multi-step Time Series forecasting, CopulaCPTS. We prove that CopulaCPTS has finite-sample validity guarantee. On four synthetic and real-world multivariate time series datasets, we show that CopulaCPTS produces more calibrated and efficient confidence intervals for multi-step prediction tasks than existing techniques. Our code is open-sourced at https://github.com/Rose-STL-Lab/CopulaCPTS.
@@ 1 Introduction
Deep learning models are becoming widely used in high-risk settings such as healthcare and transportation. In these settings, it is important that a model produces calibrated uncertainty to reflect its own confidence. Confidence regions are a common approach to quantify prediction uncertainty (Khosravi et al., 2011). A $(1-\alpha)$-confidence region $\Gamma^{1-\alpha}$ for a random variable $y$ is valid if it contains $y$’s true value with high probability: ${\mathbb{P}}[y\in\Gamma^{1-\alpha}]\geq 1-\alpha.$ Note that one can make the confidence region infinitely large to satisfy validity. But for the confidence region to be useful, we also want to minimize its area while remaining valid; this is known as the efficiency of the region.
Conformal prediction (CP) is a powerful framework to produce confidence regions with finite-sample guarantees of validity (Vovk et al., 2005; Lei et al., 2018). Furthermore, it makes no assumptions about the underlying prediction model or the data distribution. CP’s generality, simplicity, and statistical guarantees have made it popular for many real-world applications including time series prediction (Xu & Xie, 2021), drug discovery (Eklund et al., 2015) and safe robotics (Luo et al., 2021).
@@
This paper considers the setting of multi-step time series forecasting from a set of independent sequences. Consider the problem of vehicle trajectory prediction, illustrated in Figure 1. Given a dataset of trajectories, the task is to predict a future trajectory for $k$ steps given its past trajectory of $t$ time steps. We assume that the trajectories are independent from each other. For each trajectory, these time steps are temporally dependent.
There are many real-world tasks that present the same challenges as the example above, such as EEG forecasting (each patient is independent), short-term weather forecasting (local meteorology history is independent), etc. They require predicting multiple time steps into the future, so it is desired to have a “cone of uncertainty” that covers the entire course of the forecasts. Existing CP methods for time series data either only provide coverage guarantee for individual time steps (Gibbs & Candes, 2021; Xu & Xie, 2021) or produce confidence regions that are often too inefficient to be useful, especially in long horizons or multivariate settings (Stankevičiūtė et al., 2021).
In this paper, we present a practical and effective conformal prediction algorithm for multi-step time series forecasting. We introduce CopulaCPTS, a Copula-based Conformal Prediction algorithm for multi-step Time Series forecasting. A copula is a multivariate cumulative distribution function that models the dependence between multiple random variables. By using copulas to model the uncertainty jointly over future time steps, we can shrink the confidence regions significantly while maintaining validity. Copulas have been used for conformal prediction (Messoudi et al., 2021), but they focus on multiple target prediction in non-temporal settings and did not provide a validity proof.
@@
  • We introduce CopulaCPTS, a general uncertainty quantification algorithm that can be applied to any multivariate multi-step forecaster.
  • We prove that CopulaCPTS produces valid confidence regions for the full forecast horizon.
@@
  • We extend CopulaCPTS to obtain valid confidence intervals for time series forecasts of varying lengths.
@@ 2 Related Work
Deep Uncertainty Quantification for Time-Series Forecasting.
The two major paradigms of Uncertainty Quantification (UQ) methods for deep neural networks are Bayesian and Frequentist. Bayesian approaches estimate a distribution over the model parameters given data, and then marginalize these parameters to form output distributions via Markov Chain Monte Carlo (MCMC) sampling (Welling & Teh, 2011; Neal, 2012; Chen et al., 2014) or variational inference (VI) (Graves, 2011; Kingma et al., 2015; Blundell et al., 2015; Louizos & Welling, 2017). Wang et al. (2019); Wu et al. (2021) propose Bayesian Neural Networks (BNN) for UQ of spatiotemporal forecasts. In practice, Bayesian UQ can be computationally expensive and difficult to optimize, especially for larger networks (Lakshminarayanan et al., 2017; Zadrozny & Elkan, 2001). Furthermore, Bayesian methods do not provide any finite sample coverage guarantees. Therefore, UQ for deep neural network time series forecasts often adopts approximate Bayesian inference such as MC-dropout (Gal & Ghahramani, 2016b; Gal et al., 2017).
Frequentist UQ methods emphasize robustness against variations in the data. These approaches either rely on resampling the data or learning an interval bound to encompass the dataset. For time series forecasting UQ, approaches include ensemble methods such as bootstrap (Efron & Hastie, 2016; Alaa & Van Der Schaar, 2020) and jackknife methods (Kim et al., 2020; Alaa & Van Der Schaar, 2020); interval prediction methods include interval regression through proper scoring rules (Kivaranovic et al., 2020; Wu et al., 2021), and quantile regression (Takeuchi et al., 2006), with many recent advances for time series UQ (Tagasovska & Lopez-Paz, 2019; Gasthaus et al., 2019; Park et al., 2022; Kan et al., 2022). Many of the frequentist methods produce asymptotically valid confidence regions and can be categorized as distribution-free UQ techniques as they are (1) agnostic to the underlying model and (2) agnostic to data distribution.
@@
Conformal prediction (CP) is an important member of distribution-free UQ methods; we refer readers to Angelopoulos & Bates (2021) for a comprehensive introduction and survey of CP. CP has become popular because of its simplicity, generality, theoretical soundness, and low computational cost. A key feature of CP is that under the exchangeability assumption, conformal methods guarantee validity in finite samples (Vovk et al., 2005).
Most relevant to our work is the recent endeavor in generalizing CP to time-series forecasting. According to Stankevičiūtė et al. (2021) there are two settings: data generated from (1) one single time series or (2) multiple independent time series. For the first setting, ACI (Gibbs & Candes, 2021) and EnbPI (Xu & Xie, 2021) developed CP algorithms that relax the exchangeability assumption while maintaining asymptotic validity via online learning (former) and ensembling (later); Zaffran et al. (2022) further improves online adaptation. Sousa et al. (2022) combines EnbPI with conformal quantile regression (Romano et al., 2019) to model heteroscedastic time series. However, because these algorithms operate on one single time series, the validity guarantees do not cover the full horizon, posing issues for application in high-risk settings.
We focus on the setting where data consists of many independent time series. Stankevičiūtė et al. (2021) shares the same setting as ours but provides only a univariate time series solution. We show that their method of applying Bonferroni correction produces inefficient confidence regions, especially for multidimensional data or long prediction horizons. Messoudi et al. (2021) uses a copula function for multi-target CP for non-temporal data, creating box-like regions to account for the correlations between the labels. However, they do not provide theoretical proof and empirical results have shown that are often invalid. This paper builds upon these works and presents a novel two-step algorithm with guaranteed multivariate multi-step coverage and efficient confidence regions.
@@ 3.1 Inductive Conformal Prediction (ICP)
${\mathbb{P}}[y\in\Gamma^{1-\alpha}(x)]\geq 1-\alpha$ (1)
@@ 4 Copula Conformal Prediction for Time Series (CopulaCPTS)
Denote the time series dataset of size $n$ as $\mathcal{D}=\{z^{i}=(x_{1:t}^{i},y_{1:k}^{i})\}_{i=1}^{n}$, where $x_{1:t}\in\mathbb{R}^{t\times d}$ is $t$ input steps, and $y_{1:k}\in\mathbb{R}^{k\times d}$ is $k$ prediction steps, both with dimension $d$ at each step. Each data point $z^{i}$ is sampled i.i.d. from an unknown distribution $\mathcal{Z}$. In the multi-step forecasting setting, given a confidence level $1-\alpha$, the algorithm produces $k$ confidence regions for a test input $x_{1:t}^{n+1}$, denoted as $[\Gamma^{1-\alpha}_{1},\ldots,\Gamma^{1-\alpha}_{k}]$. We say the confidence regions are valid if all time steps in the forecast are covered:
@@ 4.1 Algorithm Details
The two calibration sets allow us to prove validity for both components of our algorithm. At the cost of using a subset of the data to calibrate a copula, our approach produces provably more efficient confidence regions compared to worst-case corrections such as union bounding in Stankevičiūtė et al. (2021) which is a lower bound for copulas (Appendix B.1), and more valid regions than Messoudi et al. (2021) (table 1).
@@
For brevity, we will use ${\mathcal{S}}1=\{s^{i}\}_{z^{i}\in\mathcal{D}_{cal-1}}$ to denote the set of nonconformity scores of data in $\mathcal{D}_{cal-1}$ and ${\mathcal{S}}2=\{s^{i}\}_{z^{i}\in\mathcal{D}_{cal-2}}$ the set of nonconformity scores of data in $\mathcal{D}_{cal-2}$. Subscript $j$ will be used to index the set of specific time steps of the scores: ${\mathcal{S}}1_{j}=\{s^{i}_{j}\}_{z^{i}\in\mathcal{D}_{cal-1}}$, ${\mathcal{S}}2_{j}=\{s^{i}_{j}\}_{z^{i}\in\mathcal{D}_{cal-2}}$.
@@
We adopt the empirical copula (Ruschendorf, 1976) for modeling and proof in this work. The empirical copula is a non-parametric method of estimating marginals directly from observation, and hence does not introduce any bias. For the joint distribution of $k$ time steps, we construct $C_{\text{empirical}}:[0,1]^{k}\rightarrow[0,1]$ as Eqn 7.
@@
Here boldface $\bm{\infty}$ is a k-dimensional vector with each $\bm{\infty}_{j}=\infty$ for $j=1,\ldots,k$.
@@
Note that the ${\mathbf{u}}^{*}$ is not and does not have to be unique; any solution that satisfies the constraint in Eq. 8 will guarantee multi-step validity (Appendix A). The minimization helps with efficiency, i.e. the sharpness of the confidence regions. We use a gradient descent algorithm for the optimization in implementation (see Appendix B.2 for details, and Appendix C.5 for a study of its effectiveness). Lastly, We obtain $(s_{1}^{*},\ldots,s_{k}^{*})$ by $\hat{F}_{j}^{-1}({\mathbf{u}}^{*}_{j})$ and construct the confidence region for each time step $j\in\{1,\ldots,k\}$ as the set of all $y_{j}\in\mathbb{R}^{d}$ such that the nonconformity score is less than $s_{j}^{*}$. Algorithm 1 summarizes the CoupulaCPTS procedure.
@@ 16       $\hat{y}^{i}_{1:k}\leftarrow\hat{f}(x_{1:t}^{i})$
Algorithm 1 Copula Conformal Time Series Prediction (CopulaCPTS)
@@ 5 Experiments
We compare our model with three representative works in different paradigms of deep uncertainty quantification: the Bayesian-motivated Monte Carlo dropout RNN (MC-dropout) by Gal & Ghahramani (2016a), the frequentist blockwise jackknife RNN (BJRNN) by Alaa & Van Der Schaar (2020), a conformal forecasting RNN (CF-RNN) by Stankevičiūtė et al. (2021), and the multi-target copula algorithm that does not have the two step calibration (Copula) by Messoudi et al. (2021). We use the same underlying prediction model for post-hoc uncertainty quantification methods BJRNN, CF-RNN, and CopulaCPTS. The MC-dropout RNN is of the same architecture but is trained separately, as it requires an extra dropout step during training and inference.
@@ 5.2 Real world datasets
Comparison of models at different horizon lengths.
@@
CopulaCPTS for Autoregressive prediction.
@@ 6 Conclusion and Discussion
The limitations of our algorithm are as follows. As CopulaCPTS requires two calibration steps, it is suitable only when there are abundant data for calibration. The validity proof relies on using the empirical copula, so it does not apply to other learnable copula classes. Future work includes (1) improving the autoregressive extension of CopulaCPTS, to achieve coverage over the whole horizon, and (2) developing online settings of CopulaCPTS for decision making.
Acknowledgement
This work was supported in part by Army-ECASE award W911NF-23-1-0231, the U.S. Department Of Energy, Office of Science under #DE-SC0022255, IARPA HAYSTAC Program, CDC-RFA-FT-23-0069, NSF Grants #2205093, #2146343, and #2134274.
We would like to thank Bo Zhao for her helpful comments on the paper.
@@ Acknowledgement
Let ${\mathcal{S}}1=\{s^{i}\}_{z^{i}\in\mathcal{D}_{cal-1}}$ be the set of nonconformity scores of data in $\mathcal{D}_{cal-1}$ and ${\mathcal{S}}2=\{s^{i}\}_{z^{i}\in\mathcal{D}_{cal-2}}$ the set of nonconformity scores of data in $\mathcal{D}_{cal-2}$. Subscript $j$ will be used to index the set of specific time steps of the scores: ${\mathcal{S}}1_{j}=\{s^{i}_{j}\}_{z^{i}\in\mathcal{D}_{cal-1}}$, ${\mathcal{S}}2_{j}=\{s^{i}_{j}\}_{z^{i}\in\mathcal{D}_{cal-2}}$.
@@
Each ${\mathbf{u}}^{i}$ can be seen as a multivariate nonconformity score for data sample $z^{i}$. We will now illustrate that an empirical copula on $\mathcal{U}$ is a rank statistic, and hence we can apply the proof of conformal prediction to prove a finite sample validity guarantee.
@@
Next, we define an empirical multivariate quantile function for $\mathcal{U}$, a set of $k$-dimensional vectors, based on the partial order defined in Eqn 11.  [^2] There exist other definitions of multivariate quantiles, but they cannot be used in place of our definition in this proof. We have chosen this form because it connects directly to the empirical copula.
@@
  • The uncertain regions are constructed as (Algorithm 1, line 17):